# On the Utility of Equivariance and Symmetry Breaking in Deep Learning Architectures on Point Clouds

### v1

@@ 4.3 Human motion prediction task
In table 3, we evaluate our models on the CMU Human Motion Capture dataset (Gross & Shi, 2001), consisting of 31 equally connected nodes, each representing a specific position on the human body during walking. Given node positions at a random frame, the objective is to predict node positions after 30 timesteps. As per Huang et al. (2022) we use the data of the 35th human subject for the experiment. We compare our models to NRI (Kipf et al., 2018), EGNN (Satorras et al., 2021), CEGNN (Ruhe et al., 2023) and CSMPN Liu et al. (2024) and show improved performance. We demonstrate that equivariant architectures enable better performance in motion prediction. See Fig:5 in Appendix.
Rapidash with internal $\mathrm{SE}(3)$ Equivariance Constraint
1 $\mathbb{R}^{3}$ $\mathrm{SE}(3)$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}>100}>100\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}>100}>100\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\sim 1\;\;}\sim 20\;\;\end{matrix}$
2 $\mathbb{R}^{3}$ $T_{3}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}5.44_{\pm{0.12}}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{4.81}}_{\pm{0.13}}\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}24.45_{\pm{0.66}}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{25.66}}_{\pm{1.53}}\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\sim 1\;\;}\sim 20\;\;\end{matrix}$
3 $\mathbb{R}^{3}$ none $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}6.88}5.93\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}>100}84.54\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\sim 1\;\;}\sim 20\;\;\end{matrix}$
4 $\mathbb{R}^{3}$ none $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}5.93}5.53\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}>100}>100\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\sim 1\;\;}\sim 19\;\;\end{matrix}$
5 $\mathbb{R}^{3}\times S^{2}$ $\mathrm{SE}(3)$ 6.17 8.71 $\sim 4\;\;$
@@
➟  7 $\mathbb{R}^{3}\times S^{2}$ $\mathrm{SE}(3)$ 4.95 6.71 $\sim 4\;\;$
8 $\mathbb{R}^{3}\times S^{2}$ $\mathrm{SO}(3)$ 7.26 10.03 $\sim 4\;\;$
9 $\mathbb{R}^{3}\times S^{2}$ $\mathrm{SO}(3)$ 4.95 7.57 $\sim 4\;\;$
10 $\mathbb{R}^{3}\times S^{2}$ $\mathrm{SO}(3)$ 5.31 7.99 $\sim 4\;\;$
11 $\mathbb{R}^{3}\times S^{2}$ $\mathrm{SO}(3)$ 5.14 7.50 $\sim 4\;\;$
[…]
13 $\mathbb{R}^{3}\times S^{2}$ none 5.29 60.82 $\sim 4\;\;$
14 $\mathbb{R}^{3}\times S^{2}$ none 5.31 65.80 $\sim 4\;\;$
Rapidash with Internal $T_{3}$ Equivariance Constraint
15 $\mathbb{R}^{3}$ $T_{3}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}6.53}5.99\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}43.80}>100\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\sim 1\;\;}\sim 19\end{matrix}$
16 $\mathbb{R}^{3}$ $T_{3}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{5.3}}_{\pm{0.04}}}24.32_{\pm{1.97}}\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{40.69}}_{\pm{0.87}}}>100_{\pm 6.65}\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\sim 1\;\;}\sim 19\end{matrix}$
17 $\mathbb{R}^{3}$ none $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}6.03}6.82\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}77.78}69.42\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\sim 1\;\;}\sim 19\end{matrix}$
18 $\mathbb{R}^{3}$ none $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}5.49}5.54\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}66.47}76.12\end{matrix}$ $\begin{matrix}{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\sim 1\;\;}\sim 19\end{matrix}$
Reference methods from literature
TFN - - - - - - - 66.90 - -
EGNN - - - - - - - 31.70 - -
CGENN - - - - - - - 9.41 - -
CSMPN - - - - - - - 7.55 - -
@@ 5 Results
Hypothesis 1 is accepted: Equivariant methods are more data-efficient then their non-equivariant counter part. The hypothesis is tested on Shapenet segmentation for various dataset size regimes as shown in Fig. 3 and Tab. 5 in the Appendix.
@@
Hypothesis 2 is accepted: The advantage of equivariant models gets more pronounced with task complexity. Although, on all experiments the equivariant models have an edge over non-equivariant models, this improvement is more pronounced in complex and equivariant tasks, such as QM9 regression and generation (Tab. 2), Shapenet generation (Tab. 1) and motion prediction (Tab. 3), in which we note that both Shapenet generation and CMU motion prediction do not require equivariance, but are complex geometric tasks.
@@
Hypothesis 3 is rejected: Increasing model capacity by increasing the number of channels does not lead to a performance gain as seen by the group convolutional methods. The performance of low vs high capacity (compared gray vs regular numbers in the tables) is often negligible. The results suggest that all of the models are already maxed-out in net capacity to reach optimal performance.
@@
Hypothesis 4 is accepted: For equivariant tasks, having less constraints on the model is beneficial. See, e.g., Tab. 1, comparing models 1-4 with models 16-19, the latter being less constrained, and performing significantly better. Consider also the QM9 experiments (Tab. 2) which is an equivariant task. All models 1,2 and 5-12 are equivariant and universal approximators and should be able to solve the problem. Models 1-2 however use constrained kernels (symmetric, distance-based) while the group convolutional models are not. The latter largely outperforms the former.
@@
Hypothesis 5 is accepted: Providing explicit geometric information can improve performance, even when breaking symmetry. This is, e.g., observed in the tables by comparing the group convolutional models that take either coordinates as scalars (breaks equivariance) or as vectors (maintains equvariance) against those models that do not take any coordinate information as input.
6 Discussion and Conclusion
@@
Discussion Beyond hypothesis insights, our basic convolutional architecture (Rapidash) performs exceptionally well on all considered tasks, often reaching state-of-the-art (SotA) performance, s.a. on ShapeNet/QM9 generation and CMU motion prediction. Surprisingly, even non-equivariant methods perform well on equivariant tasks, though not at SotA level and without equivariant generalization guarantees. These results support the "convolution is all you need" idea (Cohen et al., 2019; Bekkers et al., 2024), contrasting the popular "attention-is-all-you-need" claim (Vaswani, 2017). The generalizability of the presented hypotheses remains to be tested outside of the class of convolutional architectures, to which Rapidash belongs. Conclusion In this work, we conducted a comprehensive evaluation of equivariant and non-equivariant models across various tasks of varying geometric complexities. In the debate of equivariant vs non-equivariant models, we provide strong evidence in favor of introducing structure over mere scaling. Specifically, we find that equivariant models (1) are more data efficient, (2) increasingly outpace non-equivariant models when increasing task geometric complexity, and (3) see performance benefits from targeted symmetry breaking.
@@ A.1 Mathematical Prerequisites and Notations
Homogeneous Spaces. A group can act on spaces other than itself via a group action ${\mathcal{T}}:G\times X\rightarrow X$, where $X$ is the space on which $G$ acts. For simplicity, the action of $g\in G$ on $x\in X$ is denoted as $g\,x$. Such a transformation is called a group action if it is homomorphic to $G$ and its group product. That is, it follows the group structure: $(g\,g^{\prime})\,x{=}g\,(g^{\prime}\,x)\ \forall g,g^{\prime}\in G,x\in X$, and $e\,x{=}x$. For example, consider the space of 3D positions ${X}=\mathbb{R}^{3}$, e.g., atomic coordinates, acted upon by the group $G{=}\mathrm{SE}(3)$. A position $\mathbf{p}\in\mathbb{R}^{3}$ is roto-translated by the action of an element $(\mathbf{x},\mathbf{R})\in SE(3)$ as $(\mathbf{x},\mathbf{R})\,\mathbf{p}{=}\mathbf{R}\,\mathbf{p}+\mathbf{x}$.
@@ A.2 Rethinking Equivariant Message Passing as group convolution
Defining $a_{ij}=g_{x_{i}}^{-1}x_{j}$ as the edge feature between a pair of nodes, (1) the message function is $\phi_{m}(f_{i},f_{j},a_{ij})=k(g_{x_{i}}^{-1}x_{j})f_{j}$, (2) the permutation invariant aggregation function is the sum, and (3) the update map is $\phi_{u}(f_{i},m_{i})=m_{i}$. Indeed, the message function is a linear map which is determined by the attribute $a_{ij}$, in this case, determined by the relative pose between two nodes. This dependence on only relative pose underlies both the equivariance and weight sharing of equivariant message passing.
@@ A.3.4 Relationship to Our Work
A.4 Additional Results and Discussion
Modelnet40C dataset We present Modelnet40 classification results in Tab. 4.

### v2

@@ 4.2 Human motion prediction task
In table 1, we evaluate our models on the CMU Human Motion Capture dataset (Gross & Shi, 2001). We compare our models to NRI (Kipf et al., 2018), EGNN (Satorras et al., 2021), CEGNN (Ruhe et al., 2023) and CSMPN (Liu et al., 2024) and show improved performance. We demonstrate that equivariant architectures enable better performance in motion prediction.
4.3 Molecular property prediction & generation task
For predicting molecular properties and generating molecules, we use QM9 (Ramakrishnan R., 2014), We evaluate the prediction of molecular properties using the MAE metric and compare these with EGNN (Satorras et al., 2021), Dimenet++ (Gasteiger et al., 2022) and SE(3)- Transformer (Fuchs et al., 2020). For details see App. A.6.
@@ 5 Results
Hypothesis 3.1 is accepted: Equivariant methods are more data-efficient then their non-equivariant counter part. The hypothesis is tested on Shapenet segmentation for various dataset size regimes as shown in Fig. 3 and Tab. 6.
@@
Hypothesis 3.1 is accepted: The advantage of equivariant models gets more pronounced with task complexity. However, on all experiments, the equivariant models have an edge over non-equivariant models, this improvement is more pronounced in complex and equivariant tasks, such as QM9 regression and generation (Tab. 2), Shapenet generation and motion prediction (Tab. 1), in which we note that both Shapenet generation and CMU motion prediction do not require equivariance.
@@
Hypothesis 3.1 is rejected: Increasing model capacity by increasing the number of channels does not lead to a performance gain as seen by the group convolutional methods. The performance of low vs high capacity (compared to gray vs regular numbers in the tables (3 and 4) is often negligible. The results suggest that all of the models are already maxed out in net capacity to reach optimal performance.
@@
Hypothesis 3.1 is accepted: For equivariant tasks, having less constraints on the model is beneficial. See, e.g., Tab. 1, comparing models 1-4 with models 16-19, the latter being less constrained, and performing significantly better. Consider also the QM9 experiments (Tab. 2) which is an equivariant task. All models 1-2 and 5-12 are equivariant and universal approximators and should be able to solve the problem. Models 1-2 however use constrained kernels (symmetric, distance-based) while the group convolutional models do not. The latter largely outperforms the former.
@@
Hypothesis 3.1 is accepted: Providing explicit geometric information can improve performance, even when breaking symmetry. This is, e.g., observed in the tables by comparing the group convolutional models that take either coordinates as scalars (breaks equivariance) or as vectors (maintains equivariance) against those models that do not take any coordinate information as input.
Discussion and Conclusion
@@
Beyond hypothesis insights, our simple group convolutional-based architecture (Rapidash) performs exceptionally well on all considered tasks, often reaching state-of-the-art (SotA) performance, s.a. on QM9 generation and CMU motion prediction. Surprisingly, even non-equivariant methods perform well on equivariant tasks, though not at the SotA level and without equivariant generalization guarantees although with somewhat more complex architecture. These results support the ”convolution is all you need” idea (Cohen et al., 2019; Bekkers et al., 2024), contrasting the popular ”attention-is-all-you-need” claim (Vaswani, 2017). The generalizability of the presented hypotheses remains to be tested outside of the class of convolutional architectures.
In this work, we conducted a comprehensive evaluation of equivariant and non-equivariant models across various tasks of varying geometric complexities. In the debate of equivariant vs non-equivariant models, we provide strong evidence in favor of introducing structure over mere scaling. Specifically, we find that equivariant models (1) are more data efficient, (2) increasingly outpace non-equivariant models when increasing task geometric complexity, and (3) see performance benefits from targeted symmetry breaking.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.1 Mathematical Prerequisites and Notations
Homogeneous Spaces. A group can act on spaces other than itself via a group action $gT:G\times X\rightarrow X$, where $X$ is the space on which $G$ acts. For simplicity, the action of $g\in G$ on $x\in X$ is denoted as $g\,x$. Such a transformation is called a group action if it is homomorphic to $G$ and its group product. That is, it follows the group structure: $(g\,g^{\prime})\,x{=}g\,(g^{\prime}\,x)\ \forall g,g^{\prime}\in G,x\in X$, and $e\,x{=}x$. For example, consider the space of 3D positions ${X}=\mathbb{R}^{3}$, e.g., atomic coordinates, acted upon by the group $G{=}\mathrm{SE}(3)$. A position $\mathbf{p}\in\mathbb{R}^{3}$ is roto-translated by the action of an element $(\mathbf{x},\mathbf{R})\in SE(3)$ as $(\mathbf{x},\mathbf{R})\,\mathbf{p}{=}\mathbf{R}\,\mathbf{p}+\mathbf{x}$.
@@ A.2 Rethinking Equivariant Message Passing as group convolution
Defining $a_{ij}=g_{x_{i}}^{-1}x_{j}$ as the edge feature between a pair of nodes, (1) the message function is $\phi_{m}(f_{i},f_{j},a_{ij})=k(g_{x_{i}}^{-1}x_{j})f_{j}$, (2) the permutation invariant aggregation function is the sum, and (3) the update map is $\phi_{u}(f_{i},m_{i})=m_{i}$. Indeed, the message function is a linear map that is determined by the attribute $a_{ij}$, in this case, determined by the relative pose between two nodes. This dependence on only relative pose underlies both the equivariance and weight sharing of equivariant message passing.
@@ A.3.4 Relationship to Our Work
A.4 Diffusion Models
Unlike EDM (Hoogeboom et al., 2022) that uses DDPM-like diffusion model with a deterministic sampler, we use a stochastic sampler proposed in (Karras et al., 2022). The generative model used in this work, follows closing to the preconditioning diffusion model proposed in (Karras et al., 2022). We condition the diffusion model with feature scaling and noise scaling and combine outputs with skip connections, allowing for faster sampling. The sampler used in this work, implements Stochastic differential equation with a second-order connection, proposed in (Karras et al., 2022). Combining preconditioning with second-order correction sampler, the diffusion process is trained with fewer steps, resulting in fewer sampling steps.
A.5 Additional Results and Discussion
Results with extra channels In this section, we present additional results to support the hypothesis presented in the paper. In the Tab: 3 and Tab:4, we show an ablation study on ShapeNet3D dataset for part segmentation task as well as the CMU motion capture dataset for human motion prediction task. Here we present results with two different settings of hidden features C = 256 (gray) and C = 2048. The latter inflated model was trained to match the representation capacity of the rest of the models.
Rapidash with internal $\mathrm{SE}(3)$ Equivariance Constraint
[…]
1 $\mathbb{R}^{3}$ $\mathrm{SE}(3)$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}80.26_{\pm 0.06}}\\ 80.31_{\pm 0.15}\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}80.33}_{\pm 0.13}}\\ 80.20_{\pm 0.07}\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}\sim 1\;\;}\\ \sim 10\end{matrix}$
2 $\mathbb{R}^{3}$ $T_{3}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}83.95_{\pm 0.06}}\\ 83.87_{\pm 0.09}\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}52.09_{\pm 0.87}}\\ 49.63_{\pm 0.87}\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}\sim 1\;\;}\\ \sim 10\end{matrix}$
3 $\mathbb{R}^{3}$ none $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}84.23_{\pm 0.08}}\\ 84.01_{\pm 0.06}\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}34.15_{\pm 0.05}}\\ 32.22_{\pm 0.27}\end{matrix}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}65.92}_{\pm 0.98}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}\sim 1\;\;}\\ \sim 10\end{matrix}$
4 $\mathbb{R}^{3}$ none $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}84.75}_{\pm 0.02}}\\ 84.48_{\pm 0.16}\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}34.07_{\pm 0.43}}\\ 32.90_{\pm 0.47}\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}\sim 1\;\;}\\ \sim 10\end{matrix}$
Rapidash with Internal $T_{3}$ Equivariance Constraint
16 $\mathbb{R}^{3}$ $T_{3}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}85.26_{\pm 0.01}}\\ 85.38_{\pm 0.05}\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}31.41_{\pm 0.86}}\\ 32.78_{\pm 0.59}\end{matrix}$ $67.49_{\pm 0.41}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}\sim 1\;\;}\\ \sim 10\end{matrix}$
17 $\mathbb{R}^{3}$ $T_{3}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}85.82}_{\pm 0.10}}\\ 85.71_{\pm 0.17}\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}35.42}_{\pm 0.98}}\\ 32.25_{\pm 0.07}\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}\sim 1\;\;}\\ \sim 10\end{matrix}$
18 $\mathbb{R}^{3}$ none $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}85.51_{\pm 0.06}}\\ 85.52_{\pm 0.09}\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}32.79_{\pm 0.74}}\\ 31.11_{\pm 0.15}\end{matrix}$ $\textbf{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}65.06}_{\pm 1.05}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}\sim 1\;\;}\\ \sim 10\end{matrix}$
19 $\mathbb{R}^{3}$ none $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}85.66_{\pm 0.02}}\\ 85.16_{\pm 0.06}\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}33.55_{\pm 0.49}}\\ 30.62_{\pm 0.21}\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}\sim 1\;\;}\\ \sim 10\end{matrix}$
Reference methods from literature
LION - - - - - - - - - 51.85 -
PVD - - - - - - - - - 58.65 -
DPM - - - - - - - - - 62.30 -
DeltaConv - - - - - - - 86.90 - - -
PointNeXt - - - - - - - 87.00 - - -
GeomGCNN - - - - - - - 89.10 - - -
is trained with 1024 points, instead of 2048   .
Rapidash with internal $SE(3)$ Equivariance Constraint
1 $\mathbb{R}^{3}$ $\mathrm{SE}(3)$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}>100}\\ >100\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}>100}\\ >100\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}\sim 1\;\;}\\ \sim 20\;\;\end{matrix}$
2 $\mathbb{R}^{3}$ $T_{3}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}5.44_{\pm{0.12}}}\\ {\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{4.81}}_{\pm{0.13}}\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}24.45_{\pm{0.66}}}\\ {\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{25.66}}_{\pm{1.53}}\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}\sim 1\;\;}\\ \sim 20\;\;\end{matrix}$
3 $\mathbb{R}^{3}$ none $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}6.88}\\ 5.93\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}>100}\\ 84.54\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}\sim 1\;\;}\\ \sim 20\;\;\end{matrix}$
4 $\mathbb{R}^{3}$ none $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}5.93}\\ 5.53\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}>100}\\ >100\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}\sim 1\;\;}\\ \sim 19\;\;\end{matrix}$
Rapidash with Internal $T_{3}$ Equivariance Constraint
15 $\mathbb{R}^{3}$ $T_{3}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}6.53}\\ 5.99\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}43.80}\\ >100\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}\sim 1\;\;}\\ \sim 19\end{matrix}$
16 $\mathbb{R}^{3}$ $T_{3}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{5.3}}_{\pm{0.04}}}\\ 24.32_{\pm{1.97}}\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\textbf{40.69}}_{\pm{0.87}}}\\ >100_{\pm 6.65}\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}\sim 1\;\;}\\ \sim 19\end{matrix}$
17 $\mathbb{R}^{3}$ none $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}6.03}\\ 6.82\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}77.78}\\ 69.42\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}\sim 1\;\;}\\ \sim 19\end{matrix}$
18 $\mathbb{R}^{3}$ none $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}5.49}\\ 5.54\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}66.47}\\ 76.12\end{matrix}$ $\begin{matrix}{\color[rgb]{0.40,0.40,0.40}\definecolor[named]{pgfstrokecolor}{rgb}{0.40,0.40,0.40}\sim 1\;\;}\\ \sim 19\end{matrix}$
Reference methods from literature
TFN - - - - - - - 66.90 - -
EGNN - - - - - - - 31.70 - -
CGENN - - - - - - - 9.41 - -
CSMPN - - - - - - - 7.55 - -
Classification task on Modelnet40 dataset ModelNet40 dataset (Wu et al., 2015) contains 9,843 training and 2,468 testing meshed CAD models belonging to 40 categories. We present Modelnet40 classification results in Tab. 5.